OpenStudy (anonymous):

PLEASE HELP, MEDAL AWARDED Write an equation of a line that passes through the point (8, 4) and is parallel to the line y = 4x + 2. y = 4x − 28 y = 4x + 28 y = x − 2 y = x + 2

OpenStudy (anonymous):

they have the same slope since they are parallel... you can use point-slope form to find the equation...

OpenStudy (haseeb96):

first take out the slope from the given equation that is m1=4 and the line is parallel to this line so slope will be equal m1=m2 m2 =4 ookay now just point intercept form and put this slope and given point in that ( y-y1) =m (x-x1) (x1,y1)=(8,4) m=4 now this is yur job

OpenStudy (anonymous):

from slope-intercept form...\[y=mx+b\]\[y=4x+2\]the slope of the line is m=4

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

wait i can't go anywhere if y=4x+2 like where do i go now

OpenStudy (haseeb96):

A is the correct answer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you can use point-slope form for the required line since it passes thru' the given point (8,4)...\[(y-y_1)=m(x-x_1)\]

OpenStudy (haseeb96):

y-4 =4(x-8) y-4 = 4x -32 y=4x-32+4 y=4x -28 do u have any confussion /? then ask me

OpenStudy (haseeb96):

most welcome